The side of a square is 3 cm smaller than one of the sides of a rectangle and 2 cm greater than its other side. Find the side of the square, if it’s known that the area of the square is 30 cm2 less than the area of the rectangle.
Answer by ankor@dixie-net.com(22740)

The side of a square is 3 cm smaller than one of the sides of a rectangle and 2 cm greater than its other side.
Find the side of the square, if it’s known that the area of the square is 30 cm2 less than the area of the rectangle.
:
let x = the side of the square
then
(x+3) = the side of the rectangle
and
(x-2) = the other side of the rectangle
:
Rect area - square area = 30 sq/cm
(x+3)(x-2) - x^2 = 30
FOIL
x^2 + 3x - 2x - 6 - x^2 = 30
combine like terms
x - 6 = 30
x = 36 cm is the side of the square
:
see if that checks out
(39 * 34) - 36^2 =
1326 - 1296 = 30